For the Otto cycle of air shown, P1= 1 bar, T1 = 290 K, V1 = 400
cm3. If T3 = 2200 K and compression ratio is 8, find (a) the heat addition, in kJ, (b) the net work, in kJ, (c) the thermal efficiency.
The mass of the air is 0.48 g.
1
Expert's answer
2020-11-23T10:32:32-0500

(a)

"Q_1=\nu C_V(T_3-T_2)"

"\gamma=1.4"

"n=\frac{V_1}{V_2}=8"

"C_V=\frac{R}{\gamma-1}"

"\nu=\frac mM"

"p_2=p_1(\frac{V_1}{V_2})^\gamma =p_1n^\gamma"

"T_2=T_1\cdot \frac{p_2V_2}{p_1V_1}=T_1 \cdot \frac{n^\gamma}{n}=T_1n^{\gamma-1}=290\cdot8^{0.4}=666 K"

"Q_1=\frac{0.48}{29}\cdot \frac{8.31}{0.4}\cdot(2200-666)=0.527 kJ"

(b)

"Q_2=\nu C_V(T_4-T_1)"

"p_3=p_2\frac{T_3}{T_2}=p_1n^\gamma\frac{T_3}{T_2}"

"p_4=p_3(\frac{V_2}{V_1})^\gamma=p_1n^\gamma\frac{T_3}{T_2}n^{-\gamma}=p_1\frac{T_3}{T_2}"

"T_4=T_3\frac{p_4V_4}{p_3V_3}=T_3\frac{p_1\frac{T_3}{T_2}}{p_1n^\gamma\frac{T_3}{T_2}}n=T_3n^{1-\gamma}=2200\cdot 8^{-0.4}=958 K"

"Q_2=\frac{0.48}{29}\cdot \frac{8.31}{0.4}\cdot(958-290)=0.230 kJ"

(c)

"\eta=1-\frac{1}{n^{\gamma-1}}=1-\frac{1}{8^{0.4}}=56" (%)